OpenStudy (anonymous):

Please help, don't understand ! Show that [D(fg)(a)]x = (Df(a)x)(g(a)) + (Dg(a)x)(f(a)) The hint is to try x = e1 and x = e2, but idk how to go about showing that .. matrix way, or .. ?

OpenStudy (turingtest):

what does [D(fg)(a)]x represent?

OpenStudy (anonymous):

the derivative of the product of fg multiplied by a vector x

OpenStudy (turingtest):

at some point "a"? or is "a" variable?

OpenStudy (anonymous):

it is differentiable at some vector a belonging to Rn and x belongs to Rn

OpenStudy (anonymous):

On my sheet theres an arrow over the a, so I'm assuming its a vector as well

OpenStudy (turingtest):

\[[D(fg)(\vec a)]\vec x=D(f(\vec a)\vec x)g(\vec a)+D(g(\vec a)\vec x)f(\vec a)~~~~~~~~~\vec a,\vec x\in\mathbb R^n\]like so?

OpenStudy (anonymous):

yes, except on my sheet, for the last part the f(a) is before the D(g(a))x

OpenStudy (anonymous):

I'm not sure if that matters .. but for the second part of the question matrices are involved, so I'm assuming it does ..

OpenStudy (turingtest):

\[[D(fg)(\vec a)]\vec x=D(f(\vec a)\vec x)g(\vec a)+f(\vec a)D(g(\vec a)\vec x)~~~~~~~~~\vec a,\vec x\in\mathbb R^n\]

OpenStudy (anonymous):

yes, and f,g : Rn -> R3

OpenStudy (turingtest):

seems like a toughie...

OpenStudy (turingtest):

I'm not really sure how to write what happens if you plug in \(x=\vec e_1\)

OpenStudy (anonymous):

i know .. it just says consider the cases when x = e1, x = e2 ..

OpenStudy (turingtest):

I may or may not come up with something, but I'll be thinking about it...

OpenStudy (anonymous):

thank you, appreciate it.